# Solving stochastic climate-economy models: A deep least-squares Monte Carlo approach

### v1

@@ 3.3 Uncertainty quantification
Uncertainty quantification (UQ) is a research field focused on understanding how uncertainties in model inputs, parameters, and other factors propagate through models to affect their outputs. This understanding is crucial for making informed decisions based on model predictions, particularly in complex systems where such decisions can have significant consequences. A key tool in UQ are Sobol’ indices (Sobol’, 2001), which are quantitative measures used in sensitivity analysis to apportion the variance of a model output to different input variables or combinations of input variables. By identifying the most important input variables and their interactions, Sobol’ indices guide efforts to sort out the main factors which should be studied with care in complex models.
@@
First-order Sobol’ indices $S_{i}$: These indices represent the contribution of a single input variable $X_{i}$ to the output variance $\mathbb{V}(Y)$, ignoring interaction effects with other variables:
$S_{i}=\frac{\mathbb{V}_{X_{i}}(\mathbb{E}_{X_{\sim i}}[Y\,|\,X_{i}])}{\mathbb{V}(Y)},$
where $\mathbb{E}_{X_{\sim i}}[Y\,|\,X_{i}]$ denotes the conditional expectation of $Y$ given $X_{i}$ with respect to all inputs $X$ except for $X_{i}$, and $\mathbb{V}_{X_{i}}(\cdot)$ denotes the variance with respect to $X_{i}$.
@@
$S_{T_{i}}=1-\frac{\mathbb{V}_{X_{\sim i}}(\mathbb{E}_{X_{i}}[Y\,|\,X_{\sim i}])}{\mathbb{V}(Y)},$
where $\mathbb{E}_{X_{i}}[Y\,|\,X_{\sim i}]$ denotes the conditional expectation of $Y$ with respect to $X_{i}$ given all inputs $X$ except for $X_{i}$, and $\mathbb{V}_{X_{\sim i}}(\cdot)$ denotes the variance with respect to all inputs $X$ except for $X_{i}$.
@@ 3.4 Comparison with other methods
Jensen and Traeger (2014) analyze long-term economic growth uncertainty in a DICE based assessment model with an infinite-horizon. They express uncertainty in terms of stochastic shocks to the growth rate of total factor productivity. The value function is approximated by Chebyshev polynomials, and the system is solved by value function iteration. The base model has only 3 physical state variables: capital $K_{t}$, atmospheric carbon $M_{t}$, and technology level $A_{t}$.
@@
Nordhaus (2018) considers the same DICE model version as the one used in this paper. Five uncertainties are identified, the same as those explained in Subsection 2.1. These uncertainties are treated as initial parameter uncertainties. The distributions are discretized to reduce the computational burden, thereby reducing the number of possible scenarios from an uncountably infinite amount to just a few thousands. A Monte-Carlo based parameter perturbation analysis is performed, where parameters are sampled, and then the corresponding deterministic version of the DICE model is solved. In contrast to Nordhaus (2018), we don’t need to discretize the distributions, and we need to solve the model only once.
@@
Cai and Lontzek (2019) also study a stochastic version of the DICE model, extending the deterministic 6-dimensional model to a stochastic 9-dimensional model. Two additional model dimensions are due to uncertainty in the evolution of total factor productivity, and one additional dimension is due to a stochastic tipping point process. The stochastic processes are discretized, and the resulting model is solved by value function iteration, where the value function is approximated by Chebychev polynomials. The model is solved with the Blue Waters supercomputer, using 110,688 cores in parallel, with computation times of up to 8 hours. While we do not include a tipping point process in this paper, our simulation based method drastically reduces the computational burden by solving our 11-dimensional (in contrast to the 9-dimensional version of Cai and Lontzek (2019)) model formulation on a 64 core machine within around 18 hours of computation time, depending on the amount of numerical precision that is required for the solutions. Expressed in terms of pure core hours (i.e. number of cores multiplied by total computing time), this amounts to a reduction in computing time of more than $99\%$.
@@
Ikefuji et al. (2020) formulate a stochastic version of the DICE model considering one uncertainty at a time: a) uncertainty in the damage-abatement fraction, b) uncertainty in the damage parameter, c) uncertainty in the emissions-to-output ratio, and d) uncertainty in total factor productivity $A_{t}$. These uncertainties are introduced by multiplying the corresponding deterministic DICE variables by stochastic disturbances. Thus, the number of state variables is the same as in the deterministic DICE (6). To the best of our knowledge, this is the only attempt to solve a stochastic version of the DICE model by using an LSMC type approach. They use least-squares regression with polynomial basis functions to approximate the value function, i.e. in the spirit of regress later LSMC. Here, we note that their regression type Monte Carlo algorithm setup omits the integration for the conditional expectation in the Bellman equation, assuming the random disturbance is known in the transition of state variables. In principle, the standard regress later LSMC can be implemented here to handle this type of uncertainty but it will be a subject of the curse of dimensionality in the case of more than one shock.
@@
Friedl et al. (2023) present a method for solving integrated assessment models and performing uncertainty quantification. They exemplify their approach on a version of the DICE model with uncertainties in equilibrium temperature sensitivity (that contains a Bayesian learning component), and the damage function (represented by a stochastic tipping process). First, a deep neural network is trained to output, in particular, the optimal policies and value function at a given point in time, and then a Gaussian process-based model is trained to approximate quantities of interest such as the social cost of carbon in order to speed up the evaluation when calculating UQ metrics. In contrast to Friedl et al. (2023), our method approximates the conditional expectation rather than the policy functions, and then finds those by running an optimizer to solve Equation (9). Approximating $\mu_{t}$ by a regression scheme is a challenging task, since the presence of the bounds (i.e. $0\leq\mu_{t}\leq 1$) require a very careful choice of an appropriate regression scheme that can effectively interpolate the optimal policy, especially in the presence of extended periods when the policy is on the boundary. Our approach avoids this issue by finding the optimal policy through an optimizer which, once the conditional expectation has been approximated, can be performed with a high degree of numerical precision and speed. Moreover, the deep LSMC method requires performing a least-squares regression, where the loss function is the squared distance between the object of interest and the neural network prediction. This choice of loss function is significantly simpler, as it avoids the eleven individual components that enter the loss function based on an elaborate set of first-order conditions that are needed in the solution of Friedl et al. (2023). Finally, in contrast to Friedl et al. (2023), we find that there is no need to train an additional Gaussian process-based surrogate model to perform UQ for the quantities of interest (such as the social cost of carbon). Once the backward recursion (Equation (8)) has been performed, a large amount of optimal trajectories for different realizations of uncertainties can be computed easily in order to perform UQ for the quantities of interest.
@@ 4 Numerical study
In this section, we present the numerical results from applying the least-squares Monte Carlo method with transformation bias adjustment and neural network approximation of conditional expectations. For clarity, we emphazise that our state vector $X_{t}$ consists of 11 variables: the six variables from the deterministic formulation of the DICE model ($K_{t}$, $M_{t}$, $T_{t}$), the two stochastic processes $A_{t}$ and $\sigma_{t}$, as well as the three parameters discussed in Subsection 2.1 (temperature-sensitivity coefficient, damage coefficient and carbon cycle coefficient).
@@
For the backward recursion and least-squares approximation of the value function, we use $2^{23}$ sample points in the 11-dimensional state space. Figure 6 is based on $5\times 10^{5}$ forward trajectories, while the statistics reported in Table 2 are based on a sample of size $10^{6}$. To find the optimal policies in (9), we use the limited-memory Broyden–Fletcher–Goldfarb–Shanno algorithm with box constraints (L-BFGS-B). On a 64 core machine, it took between 9 hours (for $2^{22}$ samples) and 18 hours (for $2^{23}$ samples) to perform the backward recursion. Computing optimal forward trajectories then typically took around 15 minutes for $10^{5}$ trajectories, and 1 hour for $5\times 10^{5}$ trajectories.
@@
The initial year for the version of DICE model used in Nordhaus (2018) is 2015, not 2020. For illustration purposes, during calculation of the optimal forward trajectories, we made the first policy decision $(c_{0},\mu_{0})$ deterministic and equal to the optimal decision in the deterministic version of the model. This amounts to starting the forward trajectories in the year 2020 with initial values that correspond to the optimal deterministic DICE states identified in Nordhaus (2018). Moreover, the original DICE model is formulated as an infinite-horizon control problem, see Equation (1). However, our formulation of the LSMC method as discussed in Section 3 assumes a finite time horizon with $N$ time steps ($N=97$ in our case corresponding to $t=0$ being the year 2015, and $t=N$ being the year 2500). Imposing a finite time horizon corresponds to a truncation of the problem, and one needs to choose an appropriate boundary reward function $R_{N}(x)$. Similarly, as in Cai and Lontzek (2019), our terminal reward function is computed by assuming that after 2500 the policies are fixed to $\mu_{t}=1$, $c_{t}=0.78$, and that the system evolves deterministically. The reward is then equal to the discounted sum of population-weighted utility of per-capita consumption from following the fixed policies for another 100 time steps. Due to discounting and the large amount of time steps, it is assumed that a different choice of boundary reward that far ahead in the future should have a negligible impact on results for the twenty-first century.
@@
Economic capital $K_{t}$ and total factor productivity $A_{t}$ can grow quite rapidly over time, especially in scenarios where large growth in $A_{t}$ meets a low consumption rate $c_{t}$. This poses an important numerical challenge, since an appropriate domain for sampling the state variables needs to be chosen with care. A popular solution to this issue, having been applied successfully in Jensen and Traeger (2014), is to normalize economic capital as follows. First, we re-write output to express it in terms of labor-augmenting technology: $Y_{t}=A_{t}K_{t}^{\gamma}L_{t}^{1-\gamma}=K_{t}^{\gamma}(\widetilde{A}_{t}L_{t})^{1-\gamma}$, where $\widetilde{A}_{t}=A_{t}^{1/(1-\gamma)}$. Let $\widetilde{A}_{t}^{\mathrm{det}}$ denote the deterministic trajectory of $\widetilde{A}_{t}$, where $g_{A}(t)$ is fixed to be equal to the expected value. Economic capital and output are then expressed in terms of units of effective labor: $k_{t}=K_{t}/(\widetilde{A}_{t}^{\mathrm{det}}L_{t})$, and $y_{t}=Y_{t}/(\widetilde{A}_{t}^{\mathrm{det}}L_{t})$. The state variable $A_{t}$ can also be substituted by $\widetilde{A}_{t}$ and further normalized to $a_{t}=\widetilde{A}_{t}/\widetilde{A}_{t}^{\mathrm{det}}$. In our simulations, we found that these normalization steps had a favorable impact on the precision of the numerical results.
Calculating the social cost of carbon (5) requires knowledge of partial derivatives of the value function with respect to atmospheric carbon concentration and economic capital. Since we do not have an analytic representation of the value function, we follow an approximation approach that was discussed in Traeger (2014), where Chebychev polynomials were used to approximate the value function. At each time $t$, we approximate the value function $V_{t}$ by a neural network:
@@
for a suitable parameter vector $\theta$. This approach strikes a balance between numerical precision and analytical tractability, applicable even in the presence of stochasticity. Note that the idea of approximating the value function by a neural network has already been carried out in Kelly and Kolstad (2001) where, however, the neural network approximation was not used for computing the social cost of carbon.
The post-decision variables $\hat{X}_{t}$, representing the states $X_{t}$ after decision $\pi_{t}$, have the same dimension as $X_{t}$. The sampling step in Algorithm 1 requires choosing an effective sampling distribution. One standard approach would be to put a high-dimensional grid of uniformly drawn points around the deterministic DICE solution. However, in order to improve numerical precision, low-discrepancy grids are favourable in order to keep the number of sample points needed to a reasonable amount. Latin hypercube sampling offers a more favourable distribution of grid points compared to uniform sampling. We chose to use Sobol’ grid points (Sobol’, 1967), which offer even higher numerical precision compared to Latin hypercube samples. Figure 4 shows the point distribution of a uniform and of a Sobol’ grid for comparison. We found that using a low-discrepancy grid improved the numerical precision of the results.
@@
Abbildung 6: Evolution of the six most important variables over time.
@@
[…]
@@
    Productivity has a strong influence on output, but neither on damages nor on temperatures.
@@
    The carbon intensity has a completely negligible impact on the model.
@@
    The temperature-sensitivity and damage coefficients have very strong impacts on the model.
Abbildung 7: First-order (left) and total-order (right) Sobol’ indices for various model outputs with respect to uncertainty in total factor productivity (TFP), carbon intensity (SIG), temperature-sensitivity coefficient (TSC), damage coefficient (DC) and carbon cycle coefficient (CC).
@@
Tabelle 2: Statistics for major variables
Variable Mean BG Median SD IQR CV
Social cost of carbon, 2020 30.9 28.3 28.7 12.5 16.7 0.40
Temperature, 2100 (°$\degree$°C) 3.42 3.49 3.40 0.46 0.64 0.13
Carbon concentration, 2100 (ppm) 1,342 1,344 1,339 156 217 0.12
World output, 2100 (trillions, 2015\$currency-dollar\\$\$) 833.6 795.9 811.2 203.6 271.9 0.24
Emissions, 2100 14.0 13.1 12.0 13.3 23.6 0.95
Damages, 2100 (percent output) 3.0 2.9 2.9 1.0 1.4 0.34
  • SD, IQR and CV refer to standard deviation, interquartile range and coefficient of variation, respectively. BG refers to best guess, which is the value calculated along the expected trajectory, assuming that uncertainties are set to their respective means.

### v2

@@ 3.3 Uncertainty quantification
Uncertainty quantification (UQ) is a research field focused on understanding how uncertainties in model inputs, parameters, and other factors propagate through models to affect their outputs. This understanding is crucial for making informed decisions based on model predictions, particularly in complex systems where such decisions can have significant consequences. A key tool in UQ are Sobol’ indices (sobol01), which are quantitative measures used in sensitivity analysis to apportion the variance of a model output to different input variables or combinations of input variables. By identifying the most important input variables and their interactions, Sobol’ indices guide efforts to sort out the main factors which should be studied with care in complex models.
Variance-based global sensitivity analysis based on Sobol’ indices has been advocated and used in economic applications; see harenberg2019uqgsa for a general treatment and miftakhova2021gsa for an application to DICE. Related computational-economics work also emphasizes that surrogate/ML representations of value and policy functions can facilitate UQ in high-dimensional dynamic settings; see scheidegger2019mlhd.
@@
Denote by $U=(U_{1},\dots,U_{d})$ the vector of uncertain model inputs (parameters and shock drivers), assumed mutually independent, and let $Y=G(U)$ be a scalar model output of interest (e.g., $T_{2100}$ or $\mathrm{SCC}_{0}$). In our implementation, each $U_{i}$ may represent a block of inputs (e.g., the full sequence of productivity shocks over the horizon), in which case the variance decompositions below apply verbatim with $U_{i}$ interpreted as that block.
First-order Sobol’ indices $S_{i}$: These indices represent the contribution of a single input (or input block) $U_{i}$ to the output variance $\mathbb{V}(Y)$, ignoring interaction effects with other inputs:
$S_{i}=\frac{\mathbb{V}_{U_{i}}\!\big(\mathbb{E}_{U_{\sim i}}[Y\,|\,U_{i}]\big)}{\mathbb{V}(Y)},$
where $\mathbb{E}_{U_{\sim i}}[Y\,|\,U_{i}]$ denotes the conditional expectation of $Y$ given $U_{i}$ with respect to all inputs $U$ except for $U_{i}$, and $\mathbb{V}_{U_{i}}(\cdot)$ denotes the variance with respect to $U_{i}$.
@@
$S_{T_{i}}=1-\frac{\mathbb{V}_{U_{\sim i}}\!\big(\mathbb{E}_{U_{i}}[Y\,|\,U_{\sim i}]\big)}{\mathbb{V}(Y)}.$
where $\mathbb{E}_{U_{i}}[Y\,|\,U_{\sim i}]$ denotes the conditional expectation of $Y$ with respect to $U_{i}$ given all inputs $U$ except for $U_{i}$, and $\mathbb{V}_{U_{\sim i}}(\cdot)$ denotes the variance with respect to all inputs $U$ except for $U_{i}$.
@@ 3.4 Comparison with other methods
jensen2014 analyze long-term economic growth uncertainty in a DICE based assessment model with an infinite-horizon. They express uncertainty in terms of stochastic shocks to the growth rate of total factor productivity. The value function is approximated by Chebyshev polynomials, and the system is solved by value function iteration. The base model has only 3 physical state variables: capital $K_{t}$, atmospheric carbon $M_{t}$, and technology level $A_{t}$.
@@
nordhaus2018 considers the same DICE model version as the one used in this paper. Five uncertainties are identified, the same as those explained in Subsection 2.1. These uncertainties are treated as initial parameter uncertainties. The distributions are discretized to reduce the computational burden, thereby reducing the number of possible scenarios from an uncountably infinite amount to just a few thousands. A Monte-Carlo-based parameter perturbation analysis is performed, where parameters are sampled, and then the corresponding deterministic version of the DICE model is solved. In contrast to nordhaus2018, we don’t need to discretize the distributions, and we need to solve the model only once.
@@
cai2019 also study a stochastic version of the DICE model, extending the deterministic 6-dimensional model to a stochastic 9-dimensional model. Two additional model dimensions are due to uncertainty in the evolution of total factor productivity, and one additional dimension is due to a stochastic tipping point process. The stochastic processes are discretized, and the resulting model is solved by value function iteration, where the value function is approximated by Chebyshev polynomials. The model is solved with the Blue Waters supercomputer, using 110,688 cores in parallel, with computation times of up to 8 hours. While we do not include a tipping point process in this paper, our simulation based method drastically reduces the computational burden by solving our 11-dimensional (in contrast to the 9-dimensional version of cai2019) model formulation on a 64 core machine within around 18 hours of computation time, depending on the amount of numerical precision that is required for the solutions. Expressed in terms of pure core-hours (cores $\times$ wall-clock hours) - a coarse normalization and not an apples-to-apples performance metric across different model variants - this corresponds to a reduction of more than $99\%$ relative to the implementation reported in cai2019.
@@
ikefuji2020expected formulate a stochastic version of the DICE model considering one uncertainty at a time: a) uncertainty in the damage-abatement fraction, b) uncertainty in the damage parameter, c) uncertainty in the emissions-to-output ratio, and d) uncertainty in total factor productivity $A_{t}$. These uncertainties are introduced by multiplying the corresponding deterministic DICE variables by stochastic disturbances. Thus, the number of state variables is the same as in the deterministic DICE (6). To the best of our knowledge, this is the only attempt to solve a stochastic version of the DICE model by using an LSMC type approach. They use least-squares regression with polynomial basis functions to approximate the value function, i.e. in the spirit of regress later LSMC. Here, we note that their regression type Monte Carlo algorithm setup omits the integration for the conditional expectation in the Bellman equation, assuming the random disturbance is known in the transition of state variables. In principle, the standard regress later LSMC can be implemented here to handle this type of uncertainty but it will be a subject of the curse of dimensionality in the case of more than one shock.
@@
friedl2023deep present a method for solving integrated assessment models and performing uncertainty quantification. They exemplify their approach on a version of the DICE model with uncertainties in equilibrium climate sensitivity (that contains a Bayesian learning component), and the damage function (represented by a stochastic tipping process). First, a deep neural network is trained to output, in particular, the optimal policies and value function at a given point in time, and then a Gaussian process-based model is trained to approximate quantities of interest such as the social cost of carbon in order to speed up the evaluation when calculating UQ metrics. In contrast to friedl2023deep, our method approximates the conditional expectation rather than the policy functions, and then finds those by running an optimizer to solve Equation (9). Approximating $\mu_{t}$ by a regression scheme is a challenging task, since the presence of the bounds (i.e. $0\leq\mu_{t}\leq 1$) require a very careful choice of an appropriate regression scheme that can effectively interpolate the optimal policy, especially in the presence of extended periods when the policy is on the boundary. Our approach avoids this issue by computing the optimal policy via constrained numerical optimization. Once the continuation-value conditional expectation has been approximated, the per-state optimization can be solved to high numerical precision; overall wall-clock time, however, is dominated by the regression step (see Section 4 and Table 2). Moreover, the deep LSMC method requires performing a least-squares regression, where the loss function is the squared distance between the object of interest and the neural network prediction. This choice of loss function is significantly simpler, as it avoids the eleven individual components that enter the loss function based on an elaborate set of first-order conditions that are needed in the solution of friedl2023deep. Finally, in contrast to friedl2023deep, we find that there is no need to train an additional Gaussian process-based surrogate model to perform UQ for the quantities of interest (such as the social cost of carbon). Once the backward recursion (Equation (8)) has been performed, a large amount of optimal trajectories for different realizations of uncertainties can be computed easily in order to perform UQ for the quantities of interest.
Cross-paper performance comparisons are inevitably imperfect because model structures, uncertainty blocks, and numerical targets differ across contributions. To make the trade-offs assessable on a transparent yardstick, we therefore report core-hours (cores $\times$ hours) as a simple hardware-normalized summary. Table 2 summarizes the main computational features of the representative approaches discussed in this subsection.
Tabelle 2: Indicative benchmarking across representative stochastic DICE solution approaches
Reference Numerical algorithm #states Uncertainty structure Core-hours
cai2019 polynomial VFI, discretized shocks 9 productivity shock, tipping $\approx 8.9\cdot 10^{5}$
This paper deep LSMC, no discretization 11 3 parameter shocks,      2 process shocks $\approx 1.2\cdot 10^{3}$
friedl2023deep DEQN, first-order equilibrium 16 6 parameter shocks, tipping, Bayesian learning $\approx 3.2\cdot 10^{1}$
[…]
In our baseline specification, the full backward recursion (including training the deep neural network regressors at each time step) takes between 9 and 18 hours on a 64-core CPU machine, corresponding to approximately $6\cdot 10^{2}$$1.2\cdot 10^{3}$ core-hours (see Section 4 for details); subsequent uncertainty-quantification outputs are then generated via forward simulation.
The practical choice between deterministic grid/VFI approaches and simulation-based LSMC depends primarily on state dimension, the nature of uncertainty, and the intended UQ outputs. Deterministic grid methods (e.g., Chebyshev-based VFI with discretized shocks) can be attractive and highly accurate in low-dimensional settings with few shocks; however, they typically require explicit shock discretization and their computational cost grows rapidly with dimension. Simulation-based LSMC becomes preferable when (i) the state is moderately to highly multi-dimensional, (ii) uncertainty enters through multiple continuous shocks and/or parameter uncertainty without natural discretization, and (iii) one aims to produce UQ outputs (e.g., global sensitivity analysis) that require many forward evaluations. In such settings, deep neural-network regressors offer a numerically stable alternative to polynomial bases whose complexity grows quickly with dimension and interaction order (see also Section 3.2). Conversely, in small-state benchmark problems, classical VFI/projection methods or low-order polynomial LSMC may be simpler and computationally preferable.
@@ 4 Numerical study
In this section, we present the numerical results from applying the least-squares Monte Carlo method with transformation bias adjustment and neural network approximation of conditional expectations. For clarity, we emphasize that our state vector $X_{t}$ consists of 11 variables: the six variables from the deterministic formulation of the DICE model ($K_{t}$, $M_{t}$, $T_{t}$), the two stochastic processes $A_{t}$ and $\sigma_{t}$, as well as the three parameters discussed in Subsection 2.1 (temperature-sensitivity coefficient, damage coefficient and carbon cycle coefficient).
@@
For the backward recursion and least-squares approximation of the value function, we use approximately $8.4$ million ($2^{23}$) sample points in the 11-dimensional state space. Figure 6 is based on $5\times 10^{5}$ forward trajectories, while the statistics reported in Table 3 are based on a sample of size $10^{6}$. To find the optimal policies in (9), we use the limited-memory Broyden–Fletcher–Goldfarb–Shanno algorithm with box constraints (L-BFGS-B). On a 64 core machine, it took between 9 hours (for $\approx 4.2$ million, i.e. $2^{22}$, samples) and 18 hours (for $\approx 8.4$ million, i.e. $2^{23}$, samples) to perform the backward recursion. Computing optimal forward trajectories then typically took around 15 minutes for $10^{5}$ trajectories, and 1 hour for $5\times 10^{5}$ trajectories.
The numerical stability and precision of the deep LSMC recursion in this application relied primarily on three ingredients: (i) a post-decision state formulation (Eq. (18)), which reduces the effective dimension of the regression for conditional expectations; (ii) normalization of rapidly growing economic variables (capital and productivity) in effective-labor units; and (iii) low-discrepancy Sobol’ sequences for sampling the high-dimensional state space (Fig. 4). We briefly emphasize these choices here to facilitate reproducibility and to clarify which components helped reduce the computational burden in our setting.
@@
The initial year for the version of the DICE model used in nordhaus2018 is 2015 (i.e., $t=0$ under $\Delta=5$-year steps). Since our numerical discussion and UQ outputs focus on the twenty-first century, we report forward trajectories from 2020 onward (i.e., $t\geq 1$). For this purpose, we anchor the initial 2015–2020 step to the deterministic DICE benchmark and then apply the stochastic optimal policy from $t\geq 1$ under the full uncertainty specification. This convention affects only the initial step and is used solely as an initialization choice; it does not alter the deep LSMC backward recursion. Moreover, the original DICE model is formulated as an infinite-horizon control problem, see Equation (1). However, our formulation of the LSMC method as discussed in Section 3 assumes a finite time horizon with $N$ time steps ($N=97$ in our case corresponding to $t=0$ being the year 2015, and $t=N$ being the year 2500). Imposing a finite time horizon corresponds to a truncation of the problem, and one needs to choose an appropriate boundary reward function $R_{N}(x)$. This long-horizon finite truncation is standard in numerical implementations of DICE-style models: in practice (including Nordhaus-style implementations), the infinite-horizon objective is evaluated on a sufficiently long but finite horizon because discounting and convergence toward a long-run regime render further horizon extension immaterial for the welfare-relevant quantities in the twenty-first century. Similarly, as in cai2019, our terminal reward function is computed by assuming that after 2500 the policies are fixed to $\mu_{t}=1$, $c_{t}=0.78$, and that the system evolves deterministically. The reward is then equal to the discounted sum of population-weighted utility of per-capita consumption from following the fixed policies for another 100 time steps. Due to discounting and the large amount of time steps, it is assumed that a different choice of boundary reward that far ahead in the future should have a negligible impact on results for the twenty-first century.
@@
Economic capital $K_{t}$ and total factor productivity $A_{t}$ can grow quite rapidly over time, especially in scenarios where large growth in $A_{t}$ meets a low consumption rate $c_{t}$. This poses an important numerical challenge, since an appropriate domain for sampling the state variables needs to be chosen with care. A popular solution to this issue, having been applied successfully in jensen2014, is to normalize economic capital as follows. First, we re-write output to express it in terms of labor-augmenting technology: $Y_{t}=A_{t}K_{t}^{\gamma}L_{t}^{1-\gamma}=K_{t}^{\gamma}(\widetilde{A}_{t}L_{t})^{1-\gamma}$, where $\widetilde{A}_{t}=A_{t}^{1/(1-\gamma)}$. Let $\widetilde{A}_{t}^{\mathrm{det}}$ denote the deterministic trajectory of $\widetilde{A}_{t}$, where $g_{A}(t)$ is fixed to be equal to the expected value. Economic capital and output are then expressed in terms of units of effective labor: $k_{t}=K_{t}/(\widetilde{A}_{t}^{\mathrm{det}}L_{t})$, and $y_{t}=Y_{t}/(\widetilde{A}_{t}^{\mathrm{det}}L_{t})$. The state variable $A_{t}$ can also be substituted by $\widetilde{A}_{t}$ and further normalized to $a_{t}=\widetilde{A}_{t}/\widetilde{A}_{t}^{\mathrm{det}}$. In our simulations, we found that these normalization steps had a favorable impact on the precision of the numerical results.
Calculating the social cost of carbon (5) requires knowledge of partial derivatives of the value function with respect to atmospheric carbon concentration and economic capital. Since we do not have an analytic representation of the value function, we follow an approximation approach that was discussed in traeger2014, where Chebyshev polynomials were used to approximate the value function. At each time $t$, we approximate the value function $V_{t}$ by a neural network:
@@
for a suitable parameter vector $\theta$. This approach strikes a balance between numerical precision and analytical tractability, applicable even in the presence of stochasticity. Note that the idea of approximating the value function by a neural network has already been carried out in kelly2001solving where, however, the neural network approximation was not used for computing the social cost of carbon.
The post-decision variables $\hat{X}_{t}$, representing the states $X_{t}$ after decision $\pi_{t}$, have the same dimension as $X_{t}$. The sampling step in Algorithm 1 requires choosing an effective sampling distribution. One standard approach would be to put a high-dimensional grid of uniformly drawn points around the deterministic DICE solution. However, in order to improve numerical precision, low-discrepancy grids are favorable in order to keep the number of sample points needed to a reasonable amount. Latin hypercube sampling offers a more favorable distribution of grid points compared to uniform sampling. We chose to use Sobol’ grid points (sobol67), which offer even higher numerical precision compared to Latin hypercube samples. Figure 4 shows the point distribution of a uniform and of a Sobol’ grid for comparison. We found that using a low-discrepancy grid improved the numerical precision of the results.
@@
Abbildung 6: Evolution of the six most important variables over time. The deterministic DICE-2016R2 trajectories are omitted because they are nearly indistinguishable from the plotted best-guess paths and would overlap visually.
@@
Sobol’ indices quantify the contribution of each uncertain input to the variance of the output and are therefore non-directional. To make the sign of the overall impact of uncertainty transparent, Table 3 compares the best-guess value to the mean, providing a simple summary of the “uncertainty premium” for each reported quantity.
Remark 4.1.
A deterministic DICE benchmark is studied by miftakhova2021gsa, who performs a broad parametric global sensitivity analysis using a surrogate (polynomial-chaos) representation and reports Sobol’ indices for quantities such as the SCC and optimal policy. While our analysis differs in that (i) we solve a recursive stochastic-control formulation with process shocks and endogenous state-dependent policy feedback and (ii) we restrict attention to five benchmark uncertainty blocks (from nordhaus2018) rather than the full DICE parameter set, the qualitative messages for overlapping drivers are aligned: productivity uncertainty dominates output variability, whereas climate-related parameters (temperature sensitivity and damages) dominate temperature/damage variability and are important for SCC-related uncertainty. Moreover, the gap between first- and total-order indices (Figure 7) and the time variation in first-order indices (Figure 8) highlight interaction effects and time dependence that naturally arise under an endogenous optimal policy response in the stochastic formulation.
From Figure 7, output is predominantly impacted by total factor productivity, with both first-order and total-order indices around 100%, reflecting the direct Cobb–Douglas link between productivity and gross output. Productivity shocks also affect emissions and thus climate outcomes mechanically through higher output and industrial emissions; however, conditional on the benchmark uncertainty ranges and the endogenous optimal policy response, their contribution to the variance of temperature- and damage-related outputs at the reported time horizons is small relative to climate-module parameters (see the Sobol’ shares for $T^{AT}_{2100}$ and damages in Figure 7).
In contrast, within the benchmark uncertainty specification adopted from nordhaus2018, the contribution of carbon-intensity (decarbonization) uncertainty to the variance of the selected outputs is below 1% over the time horizons reported in Figure 7. This statement is meant in the variance-decomposition sense and conditional on the remaining uncertainties and on the endogenous optimal policy response; in particular, a substantial fraction of optimal-policy scenarios reaches (near) full mitigation by the end of the century, which mechanically attenuates the sensitivity of late-horizon industrial emissions to the emissions-intensity path. Accordingly, this finding should not be interpreted as a claim that recalibrated emissions-intensity pathways (e.g. SSP/RCP-based trajectories) are quantitatively unimportant; such scenario analysis is a complementary “between-scenario” layer (oneill2016scenariomip; riahi2017ssp; murakami2025scenario).
@@
    Productivity uncertainty dominates the variance of output in our benchmark set-up.
@@
    Carbon-intensity (decarbonization-rate) uncertainty contributes comparatively little to the variance of the reported quantities in this benchmark.
@@
    Climate-related parameters (climate sensitivity and the damage coefficient) account for most of the uncertainty in temperature-, damage-, and SCC-related outputs, consistent with existing DICE sensitivity analyses (see Remark 4.1 and miftakhova2021gsa).
Abbildung 7: First-order (left) and total-order (right) Sobol’ indices for various model outputs with respect to uncertainty in total factor productivity (TFP), carbon intensity (SIG), temperature-sensitivity coefficient (TSC), damage coefficient (DC) and carbon cycle coefficient ($C$).
@@
Tabelle 3: Statistics for major variables
Variable Mean BG Median SD IQR CV Mean-BG
Social cost of carbon, 2020 30.9 28.3 28.7 12.5 16.7 0.40 +9.2
[…]
Carbon concentration, 2100 (ppm) 1,342 1,344 1,339 156 217 0.12 0.1-0.1
[…]
Emissions, 2100 14.0 13.1 12.0 13.3 23.6 0.95 +6.9
Damages, 2100 (percent output) 3.0 2.9 2.9 1.0 1.4 0.34 +3.4
  • SD, IQR and CV refer to standard deviation, interquartile range and coefficient of variation, respectively. BG refers to best guess, which is the value calculated along the expected trajectory, assuming that uncertainties are set to their respective means. Mean-BG is computed as $100\times(\text{Mean}/\text{BG}-1)$ and provides the sign of the overall impact of uncertainty on each quantity (in $\%$).
These comparisons show that, in our benchmark CRRA-DICE specification, uncertainty has a modest effect on the typical (median) optimal-policy trajectory, but it can materially affect derived quantities such as the SCC through distributional asymmetry. In particular, the SCC is right-skewed at the initial date, so the mean SCC exceeds the median value. The sign and magnitude of this effect is model- and preference-specification dependent, which helps reconcile differing findings in the literature (e.g. cai2019; benmir2020green).
One reason the benchmark uncertainty effects are modest for median policy trajectories is that the standard CRRA specification links risk aversion and intertemporal substitution. More risk-sensitive recursive preferences that separate IES and risk aversion (e.g. Epstein–Zin; see epstein1989substitution; weil1990nonexpected and the IAM implementation in cai2019) can strengthen precautionary incentives and thereby amplify the effect of uncertainty on the SCC and mitigation. While we keep CRRA in the benchmark calibration to retain comparability with DICE-2016R2, our numerical framework accommodates these preference specifications with only a modification of the continuation-value recursion and regression target.